OpenStudy (anonymous):

solve for x: 3x-5+2x=25-5x

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

can you show me how you got that please

OpenStudy (anonymous):

3x-5+2x=25-5x 5x-5 =25-5x 10x-5 =25 10x =30 x =3

OpenStudy (anonymous):

Gather like terms Put x's on one side Put numbers on the other side Divide to get just x